a=299+3c
b=238+2c
how do i eliminate variables in order to continue this system of equations?
1
Expert's answer
2012-09-27T11:43:30-0400
You multiply first equation on -2 and second on 3
In result you get

-2a=-598-6c
3b=714+6c
When you add you get the result
3b-2a=116
(answer)

Other variables can't be eliminated